Solve the problem.
In a right triangle, one acute angle is 54degree more than twice the other. Find each acute angle.

Answer by Nate(3500)

The sum of the degrees in a triangle is equal to 180
One angle: x
Another angle: 2x + 54
The last angle: 90
Equation time!
(x) + (2x + 54) + (90) = 180
x + 2x + 54 + 90 = 180
3x + 144 = 180
3x = 36
x = 12
Plug:
2x + 54
2(12) + 54
24 + 54
78
One angle: x = 12
Another angle: 2x + 54 = 78
The last angle: 90